Reports until 17:23, Thursday 16 May 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 17:23, Thursday 16 May 2024 (11583)
4th super optic installed on cavity

[Torrey, Ian]

Afternoon work:

- M3:SN17 super optic is install in a flexture mount for position FCM3 in the filter cavity. https://wiki.mccullerlab.com/DCC/S2400001#chapter_dcc_s2400001_45 is updated accordingly. Both the case and flexture mount are labeled.

-I am going to keep moku 4 as the 1550 nm wavelength controller for now. Moku 1 will be 775 nm, i.e. PM driver, TRANS/REFL for 775 filter cavity

-1811 is aligned for the 775 REFL path. We dont have 775 ND filters atm. Cant do a make shift one out of a 775 PBS as we dont have any of those either. Or 775 HR with no frosted back. May need a 775 general use order to be placed.

-We are going to use a random optic we found from the previous groups leftovers to attenuate power in the mean time. It is 852 HR with no frosted back. Tested, provide sufficient power supression.

-SHG temperature controller, laser amplifier, and seeder have been shut off and unplugged in preparation of the power loss tomorrow from 4am to 8am.